Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
Direct materials 3 feet $ 5 per foot $ 15
Direct labor ? hours ? per hour ?

During March, the company purchased direct materials at a cost of $59,040, all of which were used in the production of 3,000 units of product. In addition, 4,600 direct labor-hours were worked on the product during the month. The cost of this labor time was $34,500. The following variances have been computed for the month:

Materials quantity variance $ 4,200 U
Labor spending variance $ 3,900

U

Labor efficiency variance $ 680

U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.
Answer #1
Materials Quantity variance
(AQ used - Std Qty allowed)*SR
(x - 3000*3)*5 =4200U
5x - 45000 = 4200U
x =(4200+45000)/5
9840
a) actual cost = 59040/9840
6 per foot answer
(
b) Materials price variance
(Actual rate - standard rate)*AQ purchased
(6 - 5)*9840
9840 U answer
Spending variance = 4200U+9840U
14040 U answer
2a) labor spending varian 3,900 U
labor efficiency varian 680 U
labor price variance 3220 U
labor rate variance = ( Actual rate - standard rate)*Actual hours
(34500- x*4600) = 3,220
x = (34500-3220)/4600
6.8
Standard hours 6.8 per hour
b) Standard hours =
labor efficiency variance
(Actual hours - standard hours allowed)*Standard rate
(4600 -x ) *6.8 = 680
31280 - 6.8x = 680
x                       = (31280-680)/6.8
4500
Standard hours allowed 4,500 hours
c) Standard hours allowed per unit of product
4500/3000
1.5
Standard hours allowed per unit of product 1.5 hours per unit